# Physics-informed machine learning for combustion: A review

### v1

@@ 7.2 Simulation tools (data generators)
In SciML, even with the growing number of public datasets and benchmarks, researchers still need to generate their own datasets using simulation tools, which are essential for customizing domain-specific cases, controlling data quality, and filling the inherent sparsity of public datasets. Table 7 gives a non-exhaustive list of public tools for ODE/PDE solution, fluid simulation, and combustion simulation. There are no clear boundaries between these three types of tools: Fluid simulation often falls under PDE solution, while combustion simulation often falls under either ODE solution or fluid simulation. Some ODE/PDE solution tools can be used for fluid/combustion simulation, while others cannot. For example, The CVODE 366 in SUNDIALS 367, 368 has been used in many aforementioned combustion researches 289, 276, 114, 115. The PyClaw 369, Python version of Clawpack 370, 371, has been used for simulating supersonic combustion 308. For fluid simulation, the OpenFOAM 372, 373, 374, 375 is a widely-used open-source CFD software and has many solvers for combustion, such as the chemFoam, reactingFoam, fireFoam, flameletFoam 376, and FGMFoam 377. Recently, researchers have developed tools that utilize the large language models (LLMs) to automate OpenFOAM programming, including the MetaOpenFOAM 378, 379 and OpenFOAMGPT 380, 381. Cantera 382 is an open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes, and has been used in PIML for 1D laminar flames 112. Some AI-integrated simulators are also provided in Table 7, including (1) the DeepFlame 115, a deep learning empowered open-source platform for reacting flow simulations, which coupled the three tools of OpenFOAM, Torch 383, and Cantera, and (2) differentiable solvers, including the torchdiffeq 384, $\bf{\Phi}_{\textrm{Flow}}$ 385, 386, JAX-Fluids 387, 388, Arrhenius.jl 389, and JANC 390. Table 7 also lists some tools that can be used for or are specifically designed for the DNS of fluid and combustion, including the Dedalus 391, Nek5000 392, SENGA+ 393, and AVBP 394. Note that most DNS tools for combustion are in-house codes, such as the NGA 395, S3D 396, and Athena-RFX 397, which have been used for establishing the BLASTNet 330. See more examples in this review 398 and the paper of BLASTNet 2.0 330. However, considering the extremely high cost of DNS for combustion, it is recommended to directly use the public datasets mentioned in Section 7.1 rather than generating them, especially for AI-community researchers.
@@
The tools in Table 7 have been used for generating datasets and benchmarks in Table 6. For example, the Clawpack and $\bf{\Phi}_{\textrm{Flow}}$ were used to establish the PDEBench 138, while the Clawpack and Dedalus were used to establish the Well 357. In addition to free tools, researchers can also use some widely-used commercial software for fluid/combustion simulation and dataset generation, such as the ANSYS Fluent 399, Chemkin 400, and COMSOL Multiphysics 401. For example, COMSOL Multiphysics has been used for generating data in PIML for combustion 336 and generating the datasets of PINNacle 349. An overview of both free and commercial CFD software can be found in Reference 402.
@@
SciPy 403 An open-source scientific computing library in Python language URL
SUNDIALS 367, 368 An open-source SUite of Nonlinear and DIfferential/ALgebraic Equation Solvers URL
GEKKO 404 A Python package for ML and optimization of mixed-integer and DAEs URL
@@
FEniCS 405, 406, 407 An open-source computing platform for solving PDEs based on the FEM URL
FreeFEM++ 408 A free software for solving PDEs based on the FEM URL
FiPy 409 A PDE solver using Python based on the FVM URL
Clawpack 370, 371, 369 A package that collects FVMs for hyperbolic systems of conservation laws URL
Dedalus 391 An open-source framework for solving PDEs using spectral methods URL
OpenFOAM 372, 373, 374 An open-source CFD software URL
Nek5000 392 A fast and scalable open-source CFD solver using the spectral element method URL
$\bf{\Phi}_{\textrm{Flow}}$ 385, 386 A differentiable PDE solving framework for optimization and ML applications URL
JAX-Fluids 387, 388 A differentiable CFD solver for 3D compressible single-phase and two-phase flows URL
CFD Python 410 An online learning module for CFD with Python URL
Cantera 382 An open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes URL
FlameMaster 411 An open source C++ package for 0D and 1D laminar combustion calculations URL
OpenSMOKE++ 412 A framework for simulating reacting systems with detailed kinetic mechanisms URL
SENGA+ 393 A DNS code for 3D compressible turbulent reacting flows URL
AVBP 394 A suite of CFD tools for performing LES/DNS of compressible reacting flows URL
DeepFlame 115 A deep learning empowered CFD package for reacting flows URL
Arrhenius.jl 389 A differentiable combustion simulation package URL
JANC 390 A differentiable compressible reacting flow solver URL
FDS 413 Fire dynamics simulator, a LES code for fire-driven fluid flows URL
@@ 7.3 PIML tools
There are many tools for implementing the classical PIML, i.e., PINNs, some of which are presented in Table 8. Most tools are written in Python, except for the ADCME 414 and NeuralPDE 415, which are written in Julia. Most tools can serve as solvers in which users only need to define the problem, while other tools, such as the SciANN 416 and ADCME 414, work as wrappers in which users still need to deal with most of the implementation details. Since AD has been implemented in modern AI frameworks, such as the PyTorch 417, 418, TensorFlow 419, and Keras 420, they are usually used as the backend for these PIML tools. Most PIML tools support one or two backends, except for the DeepXDE 167, which supports the PyTorch, TensorFlow, JAX 421, and PaddlePaddle 422. For more details of these PIML tools, readers can refer to these reviews 65, 77, 91, 89, where Reference 89 also listed the merits and limitations of some tools.
@@
DeepXDE 167 A library for SciML and PIML URL
PyDEns 423 A Python framework for solving ODE/PDEs using NNs URL
NeuroDiffEq 424, 425 A Python package for solving DEs with NNs URL
NVIDIA PhysicsNeMo 426 An open-source framework for building, training, and fine-tuning SciML models URL
Elvet 427 An NN-based DE and variational problem solver URL
@@
IDRLnet 429 A PINN library on top of PyTorch URL
PINA 430 A Python library designed to simplify and accelerate the development of SciML URL
SciANN 416 A Keras/TensorFlow wrapper for scientific computations and PIML using NNs URL
ADCME 414 A Julia framework to solve inverse problems involving physical simulations and NNs URL
NeuralPDE 415 A Julia package for solving PDEs using PINNs URL
Some PIML tools have been used in PIML for combustion. For example, the NVIDIA PhysicsNeMo 424, previously known as NVIDIA Modulus and SimNet 431, was used in Reference 301. The DeepXDE was used for implementing the FlamePINN-1D 112 and optimizing the extended VDP system 340. Some studies also implement the PIML for combustion directly using the modern AI frameworks, including the PyTorch 304, 302, 303, 308 and TensorFlow 306, which is more flexible but cumbersome than using PIML tools. To show how popular PIML tools can be used for combustion research, a well-commented code demo is provided in Listing 31, which is based on the DeepXDE and adapted from the codes of FlamePINN-1D. Taking the 2D combustion reacting flow as example, the demo shows how to code the Equations (3.38) to (3.42) using a PIML tool, in which the convenience of AD to compute derivatives is clearly demonstrated. To illustrate how to infer unknown parameters in inverse problems, the demo also defines an unknown correction coefficient for viscosity. It is straightforward to modify the demo for 3D combustion. For more details about the implementation of IC/BC/OCs, network, training, and evaluation, and how to code other equations in Section 3 in a differentiable way, readers can refer to the codes of FlamePINN-1D 112.
@@ 8.1.1 Limitations reinforced by combustion
PIML itself has many limitations in terms of convergence, accuracy, and generalization, while combustion, as an extremely complex multi-physical process, will pose even greater challenges for PIML. As mentioned in Section 2.3, PIML has many failure modes, such as multiscale representation 127, loss imbalance/conflict 170, mode differences in multiple subdomains 183, unreasonable sampling 169, and ill-conditioned PDEs 193. The inherent multiscale and multiphysics characteristics of combustion will certainly exacerbate these difficulties. For example, the wide range of scales involved in physical processes such as flow, chemical reaction, and diffusion in combustion reacting flows places higher demands on the multiscale representation of the ML models, while the large number of PDEs and corresponding IC/BCs poses challenges for loss balancing. Moreover, as mentioned in Section 2.3.4, most combustion PDE systems can be considered ill-conditioned. Even reaction problems that do not involve flow and diffusion have the inherent stiffness issue, as detailed in Section 4.1. In fact, the ill-conditioning issue also challenges the traditional numerical methods for combustion.
@@
To deal with these difficulties, the methods introduced in Section 2.3 should be appropriately used and improved. For example, the Fourier multiscale embedding can be used to enhance the multiscale representation 127, which has already been adopted in PIML for combustion 301. However, the introduction of hyperparameters is another minor issue, which can be resolved if the scale information is known. Moreover, there is no guarantee that this method will be effective on problems with a wider range of scales. In such cases, it is necessary to combine the domain decomposition methods, which are particularly important for large-domain and long-time problems. Domain decomposition can also facilitate parallel computation, thus accelerating the simulation. It is recommended to use knowledge-based domain decomposition if the approximate distribution characteristics of the solution are known in advance. Examples include combustors with reactant, mixing, reaction, and product regions. Adaptive domain decomposition can also be proposed for more complex problems. Unlike common domain decomposition methods, for combustion problems, each subdomain can use not only different networks but also different physical laws, which is similar to the adaptive combustion models 47. To deal with the multiphysics issue, imposing hard constraints on certain physical laws can always enhance the physical consistency and generalization ability. As for the ill-conditioning issue, some preconditioning strategies should be adopted, such as the normalization. For a deeper understanding, the ill-conditioning analysis should also be conducted for PIML solving combustion PDEs, which can be enlightened by the papers mentioned in Section 2.3.4. Although the reaction stiffness has been successfully solved in reaction ODE problems (see Section 4.1), it has not been thoroughly solved in reacting flow problems. As summarized in Table 5, when using the direct finite-rate chemistry in PIML, the most detailed reaction mechanism is only a two-step mechanism. Other studies either discard, simplify, tabulate, pre-learn, or online-learn the reaction part. Among these, only tabulating and pre-learning can represent the known detailed mechanism and may serve as feasible ways to tackle the stiffness issue in reacting flows, as long as they are differentiable. In fact, tabulating and pre-learning with NNs can be viewed as preconditioning strategies and can be categorized as building surrogate models of chemical kinetics, as mentioned in Section 4.2, not for CFD, but for PIML.
@@ 8.1.2 Discrepancies between physics, (physical) model, and data
As exemplified in Sections 4 and 5, to deal with these discrepancies between physics, model, and data, three strategies can be adopted: (1) Use reduced models, i.e., keep only the certain models and discard the uncertain parts. For example, the mass/continuity equation was adopted in most studies, as shown in Table 5, because it is certain, simple, and do not involve source terms that may need to be modeled for other equations. For the same reason, the energy and species equations are discarded in many studies. (2) Use simplified models. Examples include using RANS or LES models for turbulence and using simplified reaction mechanisms for complex reactions. Filtered PDEs have been proven to be effective for PIML to handle real-world data 204. (3) Learn the models, especially the turbulence and reaction closures, which have been illustrated in References 306, 106, 107, 307. These works demonstrate that the potential of PIML in inverse problems is not limited to field reconstruction and parameter inference, but can also assist in physical modeling. In addition to designing strategies to address these discrepancies, it is also necessary to have methods to measure these discrepancies, i.e., uncertainty quantification (UQ) 213. The methods for quantifying data and model uncertainties mentioned in Section 2.3.4 can be applied to PIML for combustion.
@@ 8.1.3 Reproducibility
Reproducibility is the foundation of academic research, ensuring the reliability, transparency, and implementability of the adopted methods in the papers, thereby guaranteeing the credibility of their conclusions. Many fields face reproducibility issues, or the so-called reproducibility crisis 432, 433, and we believe that PIML for combustion is no exception. According to Reference 434, overoptimism exists in ML for fluid-related PDEs, which is caused by weak baselines and reporting biases. The former does not apply to most studies on PIML for combustion because their goal is not to exceed a baseline model. However, reporting bias is highly likely to exist, which includes selective reporting, outcome switching, and publication bias 434. In addition to the reporting bias, the reproducibility issues can also be caused by other pitfalls, including the sensitivity to hyperparameters, ambiguity in implementation details, and inaccessibility of data and codes. As demonstrated in Tables 4 and 5, most studies on PIML for combustion do not share codes, which is very different from the open source ecosystem of the AI community.
To deal with the reproducibility issues, McGreivy et al. 434 call for bottom-up cultural changes to minimize biased reporting and top-down structural reforms to reduce perverse incentives. We briefly outline some solutions at the researcher level: (1) Provide detailed model configurations, parameter settings, and evaluation metrics. (2) Improve data accessibility by either making data available or using publicly available data, such as those in Table 6. When establishing public datasets, it is preferable to follow the principles of FAIR (Findability, Accessibility, Interoperability, and Reusability) 435. When establishing public benchmarks, some suggestions can be found in Reference 347. (3) It is preferable to provide the codes, as we have done in References 112, 313 and Listing 31. When developing public software, some suggestions can be found in Reference 436. (4) Report not only positive results but also negative results. (5) Conduct replicate experiments and report the mean and variance of the results to demonstrate that the impact of randomness is negligible. (6) Conduct hyperparameter sensitivity analysis. For hyperparameters that are highly sensitive, sufficient explanation should be provided. Finally, we emphasize that the reproducibility issues can also be alleviated by the establishment of an open, active, and rigorous community for PIML for combustion, which involves the standardization of processes, tools, and definitions, as well as the open sourcing of data, codes, and tools. There is still a long way to go in this regard.

### v2

@@ 7.2 Simulation tools (data generators)
In SciML, even with the growing number of public datasets and benchmarks, researchers still need to generate their own datasets using simulation tools, which are essential for customizing domain-specific cases, controlling data quality, and filling the inherent sparsity of public datasets. Table 7 gives a non-exhaustive list of public tools for ODE/PDE solution, fluid simulation, and combustion simulation. There are no clear boundaries between these three types of tools: Fluid simulation often falls under PDE solution, while combustion simulation often falls under either ODE solution or fluid simulation. Some ODE/PDE solution tools can be used for fluid/combustion simulation, while others cannot. For example, The CVODE 373 in SUNDIALS 374, 375 has been used in many aforementioned combustion researches 292, 279, 117, 118. The PyClaw 376, Python version of Clawpack 377, 378, has been used for simulating supersonic combustion 314. For fluid simulation, the OpenFOAM 379, 380, 381, 382 is a widely-used open-source CFD software and has many solvers for combustion, such as the chemFoam, reactingFoam, fireFoam, flameletFoam 383, and FGMFoam 384. Recently, researchers have developed tools that utilize the large language models (LLMs) to automate OpenFOAM programming, including the MetaOpenFOAM 385, 386 and OpenFOAMGPT 387, 388. Cantera 389 is an open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes, and has been used in PIML for 1D laminar flames 115. Some AI-integrated simulators are also provided in Table 7, including (1) the DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations, which coupled the three tools of OpenFOAM, Torch 390, and Cantera, and (2) differentiable solvers, including the torchdiffeq 391, $\bf{\Phi}_{\textrm{Flow}}$ 392, 393, JAX-Fluids 394, 395, Arrhenius.jl 396, and JANC 397. Table 7 also lists some tools that can be used for or are specifically designed for the DNS of fluid and combustion, including the Dedalus 398, Nek5000 399, SENGA+ 400, and AVBP 401. Note that most DNS tools for combustion are in-house codes, such as the NGA 402, S3D 403, and Athena-RFX 404, which have been used for establishing the BLASTNet 336. See more examples in this review 405 and the paper of BLASTNet 2.0 336. However, considering the extremely high cost of DNS for combustion, it is recommended to directly use the public datasets mentioned in Section 7.1 rather than generating them, especially for AI-community researchers.
@@
The tools in Table 7 have been used for generating datasets and benchmarks in Table 6. For example, the Clawpack and $\bf{\Phi}_{\textrm{Flow}}$ were used to establish the PDEBench 141, while the Clawpack and Dedalus were used to establish the Well 364. In addition to free tools, researchers can also use some widely-used commercial software for fluid/combustion simulation and dataset generation, such as the ANSYS Fluent 406, Chemkin 407, and COMSOL Multiphysics 408. For example, COMSOL Multiphysics has been used for generating data in PIML for combustion 342 and generating the datasets of PINNacle 356. An overview of both free and commercial CFD software can be found in Reference 409.
@@
SciPy 410 An open-source scientific computing library in Python language URL
SUNDIALS 374, 375 An open-source SUite of Nonlinear and DIfferential/ALgebraic Equation Solvers URL
GEKKO 411 A Python package for ML and optimization of mixed-integer and DAEs URL
@@
FEniCS 412, 413, 414 An open-source computing platform for solving PDEs based on the FEM URL
FreeFEM++ 415 A free software for solving PDEs based on the FEM URL
FiPy 416 A PDE solver using Python based on the FVM URL
Clawpack 377, 378, 376 A package that collects FVMs for hyperbolic systems of conservation laws URL
Dedalus 398 An open-source framework for solving PDEs using spectral methods URL
OpenFOAM 379, 380, 381 An open-source CFD software URL
Nek5000 399 A fast and scalable open-source CFD solver using the spectral element method URL
$\bf{\Phi}_{\textrm{Flow}}$ 392, 393 A differentiable PDE solving framework for optimization and ML applications URL
JAX-Fluids 394, 395 A differentiable CFD solver for 3D compressible single-phase and two-phase flows URL
CFD Python 417 An online learning module for CFD with Python URL
Cantera 389 An open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes URL
FlameMaster 418 An open source C++ package for 0D and 1D laminar combustion calculations URL
OpenSMOKE++ 419 A framework for simulating reacting systems with detailed kinetic mechanisms URL
SENGA+ 400 A DNS code for 3D compressible turbulent reacting flows URL
AVBP 401 A suite of CFD tools for performing LES/DNS of compressible reacting flows URL
DeepFlame 118 A deep learning empowered CFD package for reacting flows URL
Arrhenius.jl 396 A differentiable combustion simulation package URL
JANC 397 A differentiable compressible reacting flow solver URL
FDS 420 Fire dynamics simulator, a LES code for fire-driven fluid flows URL
@@ 7.3 PIML tools
There are many tools for implementing the classical PIML, i.e., PINNs, some of which are presented in Table 8. Most tools are written in Python, except for the ADCME 421 and NeuralPDE 422, which are written in Julia. Most tools can serve as solvers in which users only need to define the problem, while other tools, such as the SciANN 423 and ADCME 421, work as wrappers in which users still need to deal with most of the implementation details. Since AD has been implemented in modern AI frameworks, such as the PyTorch 424, 425, TensorFlow 426, and Keras 427, they are usually used as the backend for these PIML tools. Most PIML tools support one or two backends, except for the DeepXDE 170, which supports the PyTorch, TensorFlow, JAX 428, and PaddlePaddle 429. For more details of these PIML tools, readers can refer to these reviews 68, 80, 94, 92, where Reference 92 also listed the merits and limitations of some tools.
@@
DeepXDE 170 A library for SciML and PIML URL
PyDEns 430 A Python framework for solving ODE/PDEs using NNs URL
NeuroDiffEq 431, 432 A Python package for solving DEs with NNs URL
NVIDIA PhysicsNeMo 433 An open-source framework for building, training, and fine-tuning SciML models URL
Elvet 434 An NN-based DE and variational problem solver URL
@@
IDRLnet 436 A PINN library on top of PyTorch URL
PINA 437 A Python library designed to simplify and accelerate the development of SciML URL
SciANN 423 A Keras/TensorFlow wrapper for scientific computations and PIML using NNs URL
ADCME 421 A Julia framework to solve inverse problems involving physical simulations and NNs URL
NeuralPDE 422 A Julia package for solving PDEs using PINNs URL
Some PIML tools have been used in PIML for combustion. For example, the NVIDIA PhysicsNeMo 431, previously known as NVIDIA Modulus and SimNet 438, was used in Reference 307. The DeepXDE was used for implementing the FlamePINN-1D 115 and optimizing the extended VDP system 346. Some studies also implement the PIML for combustion directly using the modern AI frameworks, including the PyTorch 310, 308, 309, 314 and TensorFlow 312, which is more flexible but cumbersome than using PIML tools. To show how popular PIML tools can be used for combustion research, a well-commented code demo is provided in Listing 32, which is based on the DeepXDE and adapted from the codes of FlamePINN-1D. Taking the 2D combustion reacting flow as example, the demo shows how to code the Equations (3.38) to (3.42) using a PIML tool, in which the convenience of AD to compute derivatives is clearly demonstrated. To illustrate how to infer unknown parameters in inverse problems, the demo also defines an unknown correction coefficient for viscosity. It is straightforward to modify the demo for 3D combustion. For more details about the implementation of IC/BC/OCs, network, training, and evaluation, and how to code other equations in Section 3 in a differentiable way, readers can refer to the codes of FlamePINN-1D 115.
@@ 8.1.1 Limitations reinforced by combustion
PIML itself has many limitations in terms of convergence, accuracy, and generalization, while combustion, as an extremely complex multi-physical process, will pose even greater challenges for PIML. As mentioned in Section 2.3, PIML has many failure modes, such as multiscale representation 130, loss imbalance/conflict 173, mode differences in multiple subdomains 186, unreasonable sampling 172, and ill-conditioned PDEs 196. The inherent multiscale and multiphysics characteristics of combustion will certainly exacerbate these difficulties. For example, the wide range of scales involved in physical processes such as flow, chemical reaction, and diffusion in combustion reacting flows places higher demands on the multiscale representation of the ML models, while the large number of PDEs and corresponding IC/BCs poses challenges for loss balancing. Moreover, as mentioned in Section 2.3.4, most combustion PDE systems can be considered ill-conditioned. Even reaction problems that do not involve flow and diffusion have the inherent stiffness issue, as detailed in Section 4.1. In fact, the ill-conditioning issue also challenges the traditional numerical methods for combustion.
@@
To deal with these difficulties, the methods introduced in Section 2.3 should be appropriately used and improved. For example, the Fourier multiscale embedding can be used to enhance the multiscale representation 130, which has already been adopted in PIML for combustion 307. However, the introduction of hyperparameters is another minor issue, which can be resolved if the scale information is known. Moreover, there is no guarantee that this method will be effective on problems with a wider range of scales. In such cases, it is necessary to combine the domain decomposition methods, which are particularly important for large-domain and long-time problems. Domain decomposition can also facilitate parallel computation, thus accelerating the simulation. It is recommended to use knowledge-based domain decomposition if the approximate distribution characteristics of the solution are known in advance. Examples include combustors with reactant, mixing, reaction, and product regions. Adaptive domain decomposition can also be proposed for more complex problems. Unlike common domain decomposition methods, for combustion problems, each subdomain can use not only different networks but also different physical laws, which is similar to the adaptive combustion models 50. To deal with the multiphysics issue, imposing hard constraints on certain physical laws can always enhance the physical consistency and generalization ability. As for the ill-conditioning issue, some preconditioning strategies should be adopted, such as the normalization. For a deeper understanding, the ill-conditioning analysis should also be conducted for PIML solving combustion PDEs, which can be enlightened by the papers mentioned in Section 2.3.4. Although the reaction stiffness has been successfully solved in reaction ODE problems (see Section 4.1), it has not been thoroughly solved in reacting flow problems. As summarized in Table 5, when using the direct finite-rate chemistry in PIML, the most detailed reaction mechanism is only a two-step mechanism. Other studies either discard, simplify, tabulate, pre-learn, or online-learn the reaction part. Among these, only tabulating and pre-learning can represent the known detailed mechanism and may serve as feasible ways to tackle the stiffness issue in reacting flows, as long as they are differentiable. In fact, tabulating and pre-learning with NNs can be viewed as preconditioning strategies and can be categorized as building surrogate models of chemical kinetics, as mentioned in Section 4.2, not for CFD, but for PIML.
@@ 8.1.2 Discrepancies between physics, (physical) model, and data
As exemplified in Sections 4 and 5, to deal with these discrepancies between physics, model, and data, three strategies can be adopted: (1) Use reduced models, i.e., keep only the certain models and discard the uncertain parts. For example, the mass/continuity equation was adopted in most studies, as shown in Table 5, because it is certain, simple, and do not involve source terms that may need to be modeled for other equations. For the same reason, the energy and species equations are discarded in many studies. (2) Use simplified models. Examples include using RANS or LES models for turbulence and using simplified reaction mechanisms for complex reactions. Filtered PDEs have been proven to be effective for PIML to handle real-world data 207. (3) Learn the models, especially the turbulence and reaction closures, which have been illustrated in References 312, 109, 110, 313. These works demonstrate that the potential of PIML in inverse problems is not limited to field reconstruction and parameter inference, but can also assist in physical modeling. In addition to designing strategies to address these discrepancies, it is also necessary to have methods to measure these discrepancies, i.e., uncertainty quantification (UQ) 216, 47. The methods for quantifying data and model uncertainties mentioned in Section 2.3.4 can be applied to PIML for combustion.
@@ 8.1.3 Reproducibility
Reproducibility is the foundation of academic research, ensuring the reliability, transparency, and implementability of the adopted methods in the papers, thereby guaranteeing the credibility of their conclusions. Many fields face reproducibility issues, or the so-called reproducibility crisis 439, 440, and we believe that PIML for combustion is no exception. According to Reference 441, overoptimism exists in ML for fluid-related PDEs, which is caused by weak baselines and reporting biases. The former does not apply to most studies on PIML for combustion because their goal is not to exceed a baseline model. However, reporting bias is highly likely to exist, which includes selective reporting, outcome switching, and publication bias 441. In addition to the reporting bias, the reproducibility issues can also be caused by other pitfalls, including the sensitivity to hyperparameters, ambiguity in implementation details, and inaccessibility of data and codes. As demonstrated in Tables 4 and 5, most studies on PIML for combustion do not share codes, which is very different from the open source ecosystem of the AI community.
To deal with the reproducibility issues, McGreivy et al. 441 call for bottom-up cultural changes to minimize biased reporting and top-down structural reforms to reduce perverse incentives. We briefly outline some solutions at the researcher level: (1) Provide detailed model configurations, parameter settings, and evaluation metrics. (2) Improve data accessibility by either making data available or using publicly available data, such as those in Table 6. When establishing public datasets, it is preferable to follow the principles of FAIR (Findability, Accessibility, Interoperability, and Reusability) 442. When establishing public benchmarks, some suggestions can be found in Reference 354. (3) It is preferable to provide the codes, as we have done in References 115, 319 and Listing 32. When developing public software, some suggestions can be found in Reference 443. (4) Report not only positive results but also negative results. (5) Conduct replicate experiments and report the mean and variance of the results to demonstrate that the impact of randomness is negligible. (6) Conduct hyperparameter sensitivity analysis. For hyperparameters that are highly sensitive, sufficient explanation should be provided. Finally, we emphasize that the reproducibility issues can also be alleviated by the establishment of an open, active, and rigorous community for PIML for combustion, which involves the standardization of processes, tools, and definitions, as well as the open sourcing of data, codes, and tools. There is still a long way to go in this regard.